OpenStudy (anonymous):

3x + y = 14 –3x + 2y = 1 ???????

OpenStudy (mrdoe):

use elimination and you left with just y in one step

OpenStudy (mrdoe):

your*

hartnn (hartnn):

Add the 2 equations, what you get ?

OpenStudy (anonymous):

Yeah I'm confuse I was checking my work idk if I'm right/: @hartnn

OpenStudy (anonymous):

confused*

hartnn (hartnn):

then show your work, we'll verify.

OpenStudy (anonymous):

3x + y = 11 –3x + 2y = 1 3y = 12 y = 4

OpenStudy (anonymous):

???

hartnn (hartnn):

3x+y = 11 or 3x+y=14 ??

OpenStudy (anonymous):

3x+y=14

hartnn (hartnn):

3x + y = 14 –3x + 2y = 1 3y = ...? y = ....?

OpenStudy (anonymous):

15 and 3?

hartnn (hartnn):

y=5

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

Oh! Got it!

hartnn (hartnn):

y=5 3x+5=14 3x=14-5=9 3x=9 x=3.

OpenStudy (mrdoe):

3x+y=14 -3x+2y=1 add them to get 3y=15, y=5, than plug in y to get 3x+5=14= 3x=9, x=3